F(x)=-3x^2+5x
a)is the point (-1,2) on the graph of f? (Yes)
b)if x=-2, wat is f(x)? what point is on the graph of f? (-22,0)?
f(x)= -3(-2)^2+5(-2)
-12-10
=-22?
c)if f(x)=-2, what is x? What point(s) are on the graph of x? (HELP)
f(x)= -2=-3^2+5x
d)what is the domain of f? (HELP)
e)list the intercepts, if any, of the graph of f.(HELP)
f)list the y intercept, if there is one, of the graph f.(HELP)
I've tried B and got x=-22, but Im stuck mainly on C and any help on the rest would be great!
Found 2 solutions by richwmiller, stanbon: Answer by richwmiller(17219) 
 f(x)= -2=-3^2+5x
not -2=-3^2+5x
but -2=-3x^2+5x
0=-3x^2+5x+2
we want factors of -6 that add up to 5
6 and -1
0=-3x^2+6x-x+2
0=-3x*(x-2)-1*(x-2)
0=-1*(3x+1)*(x-2)
Answer by stanbon(75887) 
 F(x)=-3x^2+5x
a)is the point (-1,2) on the graph of f?
f(-1) = -3(-1)^2 + 5(-1) = -3-5 = -8
Ans: No.
---------------
b)if x=-2, what is f(x)?
f(-2) = -3(4)+5(-2) = -22
---
what point is on the graph of f? (-2,-22)
---------------
c)if f(x)=-2, what is x?
-2 = -3x^2+5x
3x^2-5x-2 = 0
x = [5 +- sqrt(25-4*3*-2)]/6
x = [5+-7]/6
x = 2 or x = -1/3
Points: (2,-2) and (-1/3,-2)
=========================================
d)what is the domain of f(x) = -3x^2+5x?
Domain: All Real Numbers
---------
e)list the intercepts, if any, of the graph of f.
x-intercept= ?
Let y = 0 and solve for "x":
-3x^2+5x = 0
x(3x+5) = 0
x = 0 or x = -5/3
x-intercepts: (0,0),(-5/3,0)
=========================================
f)list the y intercept, if there is one, of the graph f.
Let x = 0, then -3*0 + 5*0 = 0
y-intercept : (0,0)
=========================================
Cheers,
Stan H.
=========================================
| |
|
| |